show me how to get the answer for 1/5 +1/8 +1/10 =
Found 2 solutions by rajagopalan, richwmiller:
Answer by rajagopalan(174)

answer for 1/5 +1/8 +1/10
LCM of 5, 8, 10 = 40
so 1/5=8/40
1/8=5/40
1/10=4/40
adding the numerators 8+5+4=17
Answer=17/40
Good day.

Answer by richwmiller(17219)

The lcd is 40
8/40+5/40+4/40=17/40